# Experimental Study of Underwater Acoustic Reconfigurable Intelligent Surfaces with In-Phase and Quadrature Modulation

### v1

@@ Abstract
This paper presents an underwater acoustic reconfigurable intelligent surfaces (UA-RIS) designed for long-range, high-speed, and environmentally friendly communication in oceanic environments. The proposed UA-RIS comprises multiple pairs of acoustic reflectors that utilize in-phase and quadrature (IQ) modulation to flexibly control the amplitude and phase of reflected waves. This capability enables precise beam steering to enhance or attenuate sound levels in specific directions. A prototype UA-RIS with 4$\times$6 acoustic reflection units is constructed and tested in both tank and lake environments to evaluate performance. The experimental results indicate that the prototype is capable of effectively pointing reflected waves to targeted directions while minimizing side lobes using passive IQ modulation. Field tests reveal that deploying the UA-RIS on the sender side considerably extends communication ranges by 28% in deep water and 46% in shallow waters. Furthermore, with a fixed communication distance, positioning the UA-RIS at the transmitter side substantially boosts data rates, with an average increase of 63.8% and peaks up to 96%. When positioned on the receiver side, the UA-RIS can expand the communication range in shallow and deep water environments by 40.6% and 66%, respectively. Moreover, placing the UA-RIS close to the receiver enhances data rates by an average of 80.3%, reaching up to 163% under certain circumstances.
@@
Reconfigurable intelligent surface (RIS), underwater acoustic communication, in-phase and quadrature (IQ) modulation, lake tests.
@@ I Introduction
In recent years, radio frequency reconfigurable intelligent surfaces (RF-RIS) have emerged as a promising technology to enhance communication quality for 5G and beyond networks [1]. By dynamically adjusting the amplitude and phase of reflected electromagnetic waves with the load network, RIS can precisely steer beams in specific directions. This capability enhances signal strength at the receiver while reducing interference in surrounding environments, thereby improving the overall performance of wireless communication networks [2].
Although extensive research has been conducted for RIS in radio environments, bringing this cutting-edge technology into underwater environments remains a complex problem. In water, the transmission range of electromagnetic (EM) waves is severely restricted due to significant absorption attenuation. Currently, acoustic communication is still the principal method for mid-range and long-range communication in marine settings. However, the physical properties of acoustic waves are completely distinct from those of radio signals, rendering existing RF-RIS unsuitable for direct application in water. At present, only limited studies explore the design of underwater acoustic RIS (UA-RIS) [3, 4].
Our work demonstrates that UA-RIS can significantly extend communication range, enhance data rates, and optimize channel utilization for underwater acoustic networks. Nevertheless, to transition UA-RIS from theory to practical application, we must overcome a series of distinct challenges: {adjustwidth}-0.71cm0cm
@@
To address the above issues, we proposed a UA-RIS, which can automatically pair acoustic reflectors based on the angles of reflected waves. Each pair of reflectors then performs in-phase and quadrature (IQ) modulation to direct the reflected beam in specific directions. Compared to the 1-bit and 2-bit coding schemes widely used in RF-RIS for beam steering [6, 1], the IQ modulation offers significantly enhanced flexibility. Specifically, traditional multi-bit coding schemes can produce only a finite number of phase states. This limitation results in coarse control over the reflected waves, leading to less precise beam direction. Furthermore, multi-bit coding schemes typically do not allow altering the amplitude of the reflected wave, resulting in higher side lobes in the beam pattern, which can cause unexpected interference issues. In contrast, IQ modulation enables the formation of reflected waves with arbitrary amplitude and phase, supporting more sophisticated array processing methods such as minimum variance distortionless response (MVDR) and linearly constrained minimum variance (LCMV) [7]. As a result, the reflected waves can be accurately directed in the desired directions without introducing strong interferences in environments.
@@
We carried out comprehensive experiments in tank and lake environments to evaluate the performance of the proposed work across various configurations. The experimental results show that the UA-RIS can flexibly enhance or attenuate the strength of acoustic waves in specific directions. By deploying the UA-RIS with 24 reflection units at the sender side, the data rate of acoustic communication can be improved by an average of 63.8% (up to 96%), or the communication rage can be extended by 28% and 46% in deep water and shadow water environments, respectively. When positioned near the receiver, the UA-RIS extends the communication range in shallow and deep water environments by 40.6% and 66%, respectively. Additionally, situating the UA-RIS adjacent to the receiver boosts data rates by an average of 80.3%, with increases up to 163% under certain conditions.
@@
An IQ modulation scheme is developed for UA-RIS that allows to generation of reflected signals with arbitrary amplitude and phase shifts. This technique ensures precise steering of the reflected waves, directing them effectively while minimizing interference in unintended directions.
@@
Both tank and lake experiments are conducted to validate the feasibility of the UA-RIS. The results confirm that our system can significantly enhance the quality of receiving signals in underwater environments for long-range and high-speed acoustic communications.
The structure of this paper is as follows: Section II reviews related work on UA-RIS. Section III explores two applications to motivate our work. Section IV examines the benefits of using IQ modulation in UA-RIS. Section V provides a detailed description of the UA-RIS hardware design. Lastly, Section VI presents an experimental validation of the system performance.
@@ II-A RIS in Terrestrial Radio Environments
The advent of RF-RIS technology lies in the study of metamaterials, which are composed of periodic structures with scales that are smaller than the wavelength of interacting electromagnetic (EM) waves. These metamaterials can resonate with incident EM fields, consequently producing a specific EM response on the surface with a precise design of the surface architecture [8, 9].
@@ II-B RIS in Underwater Acoustic Environments
The study in [4] investigates wide-band beamforming using the UA-RIS architecture developed in [3]. This approach aims to neutralize two types of dispersion effects: element dispersion within each piezoelectric reflector and array dispersion across the acoustic RIS. By doing so, the directivity of the reflected waves becomes frequency-independent, allowing wideband signals to achieve consistent beamforming results across various frequency components. The effectiveness of the proposed architecture is assessed using COMSOL and the Bellhop simulator.
III Applications of UA-RIS
In this section, we introduce two potential applications of using UA-RIS for achieving ultra-long-distance and environmentally friendly underwater acoustic communications.
III-A Environmentally Friendly Acoustic Communication
In oceans, underwater wildlife, such as dolphins and whales, depend critically on acoustic signals for communication, hunting, and navigation [21]. As shown in Fig. 1, bottlenose dolphins use frequencies from 200 Hz to 24 kHz for whistles and 200 Hz to 150 kHz for echolocation clicks, while killer whales’ echolocation typically occurs within the 12 kHz to 25 kHz range [22]. These frequencies coincide with those used by human-made acoustic devices, potentially leading to disruptions or even damage to the auditory systems of marine life. Consequently, ensuring the coexistence of marine wildlife is an essential factor in designing next-generation underwater communication systems.
The UA-RIS technology offers a viable solution to achieve the aforementioned goal. With the assistance of RIS, acoustic modems can transmit data at a lower sound level, reducing their environmental footprint, and thereby minimizing the interference of sound devices with the acoustic channel. Furthermore, UA-RIS directs the reflected signal towards a specific direction rather than dispersing them throughout the space. This approach not only diminishes interference but also improves channel utilization in the spatial domain.
III-B Long-Range Underwater Communication
For long-range communication, traditional underwater acoustic systems usually operate at low frequencies to mitigate propagation attenuation. This approach often requires either larger acoustic transducers or increased transmission power, which in turn shortens battery life. In contrast, UA-RIS provides an alternative by allowing the use of compact, low-power transducers for effective long-range communication effectively.
To achieve the above objective, multiple UA-RIS units are strategically deployed at both the ocean’s surface and its depths. As illustrated in Fig. 2, the surface UA-RIS unit is solar-powered and mounted on a buoy, while the deep-sea UA-RIS is connected to a wirewalker [23], a fully passive device that harvests energy from the motion of ocean waves.
By meticulously designing the system, UA-RIS units precisely direct the source signal toward the intended receiver during transmission. This approach coherently combines multiple acoustic waves reflected by the UA-RIS units at the receiver side, thereby substantially boosting the signal-to-noise ratio (SNR). As a result, it facilitates extended communication ranges, even at higher frequencies, by effectively preserving signal strength over long distances.
IV IQ Modulation for UA-RIS
@@
This section explores the IQ modulation technique utilized in our UA-RIS. We begin by examining the challenge of implementing arbitrary phase shifts within UA-RIS, followed by a description of how IQ modulation addresses this issue. Subsequently, we discuss the benefits of IQ modulation, highlighting its feasibility, flexibility, and interference reduction.
IV-A Challenges of Implementing Arbitrary Phase-Shift in UA-RIS
@@
According to equation (1), an RIS with a characteristic impedance of 50 $\Omega$, typical for RF transmission lines, requires load impedances of $-121j\,\Omega$ and $-50j\,\Omega$ to generate reflected waves with the same amplitude as the incident wave but with phase shifts of $-45^{\circ}$ and $-90^{\circ}$, respectively. Assuming a signal frequency of 30 GHz is selected for mmWave communication, these load impedances correspond to capacitances of approximately 44 pF and 106 pF, respectively. Commercial varactor diodes, such as the ZC836A [24] and NTE618 [25], can continuously vary capacitance within this range by adjusting the reverse bias voltage applied to their p-n junctions, thereby enabling arbitrary phase shifts of the reflected waves.
However, the frequencies used for middle-range or long-range acoustic communications are typically low, often below 50 kHz. Under these conditions, the capacitance of the varactor diode needs to change significantly to shift the phase of reflected waves. For example, maintaining the conditions from the previous analysis but reducing the signal frequency to 50 kHz, the load network in the RIS would need to adjust the capacitance from 26 nF to 64 nF to shift the phase of the incident wave from $-45^{\circ}$ to $-90^{\circ}$. This range exceeds the capabilities of most commercial varactor diodes. Consequently, a new method is essential to implement arbitrary phase shifts for UA-RIS.
IV-B Implementation of IQ modulation
In the proposed UA-RIS, we employ IQ modulation to generate reflected waves with arbitrary amplitude and phase shift. To implement this scheme, the MCU automatically pairs acoustic reflectors according to the angles of reflected waves. Subsequently, one reflector in each pair produces an in-phase or antiphase component, shifting the phase of the reflected waves by 00° or $180$°. The other reflector generates a quadrature component, shifting the phase of the reflected waves by $90$° relative to the incident signal. By precisely controlling the amplitude of the in-phase, antiphase, and quadrature components, the UA-RIS can synthesize reflected waves with any desired phase in space, as dictated by trigonometric identities:
@@ IV-B Implementation of IQ modulation
Figure 3: IQ modulation for arbitrary phase shift.
Fig. 3 illustrates how IQ modulation operates. In the diagram, each circle represents an acoustic reflector, the size of which is one wavelength to efficiently reflect the incident waves. In the figure, the green plane indicates the direction of the incident wave, while the orange plane designates the target direction for the reflected waves.
To execute IQ modulation, only reflectors located on the wavefront of the reflected wave (as marked by the black dotted line in the figure) can be paired. For example, reflector 1 can be paired with reflector 2 or 2. With this configuration, the projection positions of the paired reflectors along the direction of the reflected waves align. Consequently, positional differences between paired reflectors do not result in additional phase shifts in the direction of the reflected wave. This setup enables the paired reflectors to function collectively as a single reflector, facilitating precise beam steering.
It is important to recognize that the phase of the incident wave received by each reflector in a pair may differ. This discrepancy arises because the projection positions of the paired reflectors along the wavefront of the incident signals do not always align, as depicted in Fig. 3, where the green dotted line (representing the wavefront of the incident wave) is not parallel to the black dotted line. Therefore, the UA-RIS must correct these phase differences to ensure that the reflected waves generated by the pair of reflectors remain orthogonal, which is crucial for effective IQ modulation.
@@
Comparing (3) with (2) reveals that $\left(A_{1}\cos\phi_{1}+A_{2}\sin\phi_{2}\right)$ is the amplitude of the in-phase or antiphase component, while $\left(A_{1}\sin\phi_{1}+A_{2}\cos\phi_{2}\right)$ is the amplitude of the quadrature components for IQ modulation.
@@
V Hardware Architecture
This section outlines the hardware design of UA-RIS, beginning with an introduction to the structure of the acoustic reflector and the circuitry designed for efficient IQ modulation. Following this, we will discuss several critical considerations essential for real-world implementation.
V-A Structure and Circuit of Acoustic Reflector
V-A1 Reflector Structure

### v2

# Experimental Study of Underwater Acoustic Reconfigurable Intelligent Surfaces with Synthetic Reflection

@@ Abstract
This paper presents an underwater acoustic reconfigurable intelligent surfaces (UA-RIS) designed for long-range, high-speed, and environmentally friendly communication in oceanic environments. The proposed UA-RIS comprises multiple pairs of acoustic reflectors that utilize a synthetic reflection scheme to flexibly control the amplitude and phase of reflected waves. This capability enables precise beam steering to enhance or attenuate sound levels in specific directions. A prototype UA-RIS with 4$\times$6 acoustic reflection units is constructed and tested in both tank and lake environments to evaluate performance. Experimental results using a continuous wave (CW) as the source signal demonstrate that the prototype is capable of effectively pointing reflected waves to targeted directions while minimizing side lobes through synthetic reflection. Field tests reveal that deploying the UA-RIS on the sender side considerably extends communication ranges by 28% in deep water and 46% in shallow waters. Furthermore, with a fixed communication distance, positioning the UA-RIS at the transmitter side substantially boosts data rates, with an average increase of 63.8% and peaks up to 96%. When positioned on the receiver side, the UA-RIS can expand the communication range in shallow and deep water environments by 40.6% and 66%, respectively. Moreover, placing the UA-RIS close to the receiver enhances data rates by an average of 80.3%, reaching up to 163% under certain circumstances.
@@
Reconfigurable intelligent surface (RIS), underwater acoustic communication, synthetic reflection, lake tests.
@@ I Introduction
In recent years, radio frequency reconfigurable intelligent surfaces (RF-RIS) have emerged as a enabling technology to enhance communication quality for 5G and beyond networks [1]. By dynamically adjusting the amplitude and phase of reflected electromagnetic waves with the load network, RIS can precisely steer beams in specific directions. This capability enhances signal strength at the receiver while reducing interference in surrounding environments, thereby improving the overall performance of wireless communication networks [2].
Extensive research has been conducted for RIS in radio environments, bringing this cutting-edge technology into underwater environments remains a complex problem. In water, the transmission range of electromagnetic (EM) waves is severely restricted due to significant absorption attenuation. Currently, acoustic communication is still the principal method for mid-range and long-range communication in marine settings. However, the physical properties of acoustic waves are completely distinct from those of radio signals, rendering existing RF-RIS unsuitable for direct application in water. At present, only limited studies explore the design of underwater acoustic RIS (UA-RIS) [3, 4].
UA-RIS can significantly extend the communication range, making it a promising solution for achieving ultra-long-range underwater communication. Traditional acoustic systems typically operate at low frequencies to reduce propagation loss, but this often requires either bulky transducers or higher transmission power, both of which can reduce battery life. In contrast, UA-RIS enables the use of compact, low-power transducers while still maintaining effective long-range performance by strategically deploying multiple intelligent surfaces at both the ocean surface and seabed. As illustrated in Fig. 1, the surface UA-RIS unit is solar-powered and mounted on a buoy, while the deep-sea UA-RIS is connected to a wirewalker [5], a fully passive device that harvests energy from the motion of ocean waves.
Through careful design, UA-RIS units can precisely direct the source signal toward the intended receiver, coherently combining reflected acoustic waves to significantly enhance the signal-to-noise ratio (SNR). This enables extended communication ranges, even at higher frequencies, by effectively preserving signal strength over long distances.
In addition to extending the communication distance, UA-RIS can reduce the source level of acoustic signals without compromising communication quality, enabling more environmentally friendly underwater communications. Marine wildlife, such as dolphins and whales, rely heavily on sound for communication, hunting, and navigation [6]. As shown in Fig. 2, bottlenose dolphins use frequencies from 200 Hz to 24 kHz for whistles and 200 Hz to 150 kHz for echolocation clicks, while killer whales’ echolocation typically occurs within the 12 kHz to 25 kHz range [7]. These ranges overlap with those of human-made acoustic systems, potentially causing interference or auditory harm.
UA-RIS technology provides a viable solution to the above challenge. By enabling acoustic modems to transmit at lower sound levels, it reduces environmental impact and minimizes interference with the acoustic channel. Additionally, UA-RIS focuses reflected signals in specific directions rather than dispersing them, thereby reducing interference and enhancing spatial channel utilization.
Although UA-RIS offers advantages in long-range and environmentally friendly communication, transitioning this technology from theory to practical application requires overcoming several key challenges:
{adjustwidth}
-0.71cm0cm
@@
To address the above issues, we proposed a new UA-RIS architecture, which automatically pair acoustic reflectors based on the angles of reflected waves. Each pair of reflectors then performs synthetic reflection scheme to direct the reflected beam toward specific directions. Compared to the 1-bit and 2-bit coding schemes widely used in RF-RIS for beam steering [9, 1], the synthetic reflection offers significantly enhanced flexibility. Specifically, traditional multi-bit coding schemes can produce only a finite number of phase states. This limitation results in coarse control over the reflected waves, leading to less precise beam direction. Furthermore, multi-bit coding schemes typically do not allow altering the amplitude of the reflected wave, resulting in higher side lobes in the beam pattern, which can cause unexpected interference issues. In contrast, synthetic reflection enables the formation of reflected waves with arbitrary amplitude and phase, supporting more sophisticated array processing methods such as minimum variance distortionless response (MVDR) and linearly constrained minimum variance (LCMV) [10]. As a result, the reflected waves can be accurately directed in the desired directions without introducing strong interferences in environments.
@@
We carried out comprehensive simulations using COMSOL Multiphysics, along with real-world experiments in both tank and lake environments, to evaluate the performance of the proposed work across various configurations. The experimental results using a CW signal as the source show that the UA-RIS can flexibly enhance or attenuate the strength of acoustic waves in specific directions. By deploying the UA-RIS with 24 reflection units at the sender side, the data rate of acoustic communication can be improved by an average of 63.8% (up to 96%), or the communication range can be extended by 28% and 46% in deep water and shadow water environments, respectively. When positioned near the receiver, the UA-RIS extends the communication range in shallow and deep water environments by 40.6% and 66%, respectively. Additionally, situating the UA-RIS adjacent to the receiver boosts data rates by an average of 80.3%, with increases up to 163% under certain conditions.
@@
An synthetic reflection scheme is developed for UA-RIS that allows to generation of reflected signals with arbitrary amplitude and phase shifts. This technique ensures precise steering of the reflected waves, directing them effectively while minimizing interference in unintended directions.
@@
Comprehensive validation of the UA-RIS was performed through tank and lake experiments as well as COMSOL Multiphysics simulations. The results demonstrate that the proposed system can significantly enhance signal reception quality in underwater environments, enabling long-range and high-speed acoustic communications.
The structure of this paper is as follows: Section II reviews related work on UA-RIS. Section III examines the benefits of using synthetic reflection in UA-RIS. Section IV provides a detailed description of the UA-RIS hardware design. Lastly, Section V presents an experimental validation of the system performance.
@@ II-A RIS in Terrestrial Radio Environments
The advent of RF-RIS technology lies in the study of metamaterials, which are composed of periodic structures with scales that are smaller than the wavelength of interacting electromagnetic (EM) waves. These metamaterials can resonate with incident EM fields, consequently producing a specific EM response on the surface with a precise design of the surface architecture [11].
@@ II-B RIS in Underwater Environments
The study in [4] investigates wide-band beamforming using the UA-RIS architecture developed in [3]. This approach aims to neutralize two types of dispersion effects: element dispersion within each piezoelectric reflector and array dispersion across the acoustic RIS. By doing so, the directivity of the reflected waves becomes frequency-independent, allowing wideband signals to achieve consistent beamforming results across various frequency components.The effectiveness of the proposed architecture is assessed using COMSOL Multiphysics and the Bellhop simulator. While the proposed UA-RIS architecture demonstrates promising performance in simulation, its practical feasibility remains to be validated through real-world experiments.
The concept of RIS has also been extended to underwater optical communications to enable non-line-of-sight (NLOS) links when direct line-of-sight (LOS) paths are obstructed by obstacles such as schools of fish, submarines, or seamounts [23, 24]. In [23], an analytical study of an RIS-assisted underwater optical communication system is presented. It models a decode-and-forward RIS relay under mixed Exponential-Gamma oceanic turbulence and Rayleigh fading, deriving closed-form expressions for end-to-end outage probability and average BER. More recently, the research in [24] tackled the challenge of “skip-zones” at the air-water interface by placing RIS at the boundary of an free-space optical communication link. By incorporating models of Gamma-Gamma atmospheric turbulence, generalized Gamma oceanic turbulence, and pointing errors, it demonstrates that RIS can bridge up to 30 meters of optical dead zone and improve BER by five orders of magnitude at 1 Gbps.
Parallel-relay designs have also been explored in [25]. This research proposed a dual-hop system with separate free-space optical communication and underwater wireless optical communication decode-and-forward relays at different depths, selecting the best branch by instantaneous SNR. Their analysis under worst-case turbulence and misalignment predicts a $\sim$3 dB gain over single-relay schemes at $10^{-4}$ outage, with free-space optical communication dominating in clear-water conditions. On the hardware front, [26] demonstrated a 16$\times$16 liquid-crystal spatial light modulator RIS with integrated aperture-averaging lens and phased-array beamformer. In lab tests at 10 Mbps over 15 m, adaptive beam steering achieved BER less than $10^{-5}$ and a fourfold SNR improvement compared to no-RIS.
Although underwater optical communication offers high data rates, its effective range is typically under 100 meters [23, 24, 25, 26]. In contrast, acoustic RIS leverages lower frequencies and longer wavelengths to achieve communication to kilometer scales even in turbid, particle-rich waters that rapidly attenuate optical beams, making UA-RIS ideal for low-power, long-range underwater communications in challenging environments.
Our work presents two key innovations that distinguish it from prior efforts on UA-RIS [4, 3]. First, we propose a novel approach to phase control in UA-RIS that addresses practical limitations associated with existing architectures. While previous studies have explored varactor-based impedance tuning, an approach well-suited to RF applications but challenging at acoustic frequencies due to the significantly larger capacitance variation required. To address this challenge, we introduce a synthetic reflection technique that achieves phase control by combining orthogonal components generated by two independently controlled reflectors. This design allows for flexible and accurate phase adjustment using readily available components. Second, whereas earlier works have primarily focused on theoretical analysis and simulation, our study includes comprehensive experimental validation of the proposed architecture. We conduct both tank experiments and open-water lake tests to evaluate the effectiveness and feasibility of the proposed architecture in real-world conditions. To the best of our knowledge, this is the first field-deployed and experimentally verified UA-RIS, providing a critical step toward practical underwater acoustic communication using intelligent surfaces.
III Synthetic Reflection for UA-RIS
@@
This section explores the synthetic reflection technique utilized in our UA-RIS. We begin by examining the challenge of implementing arbitrary phase shifts within UA-RIS, followed by a description of how synthetic reflection addresses this issue. Subsequently, we discuss the benefits of synthetic reflection, highlighting its feasibility, flexibility, and interference reduction.
III-A Challenges of Implementing Arbitrary Phase-Shift in UA-RIS
@@
According to (1), an RIS with a characteristic impedance of 50 $\Omega$, typical for RF transmission lines, requires load impedances of $-121j\,\Omega$ and $-50j\,\Omega$ to generate reflected waves with the same amplitude as the incident wave but with phase shifts of $-45^{\circ}$ and $-90^{\circ}$, respectively. Assuming a signal frequency of 30 GHz is selected for mmWave communication, these load impedances correspond to capacitances of approximately 44 fF and 106 fF, respectively. Commercial varactor diodes, such as the ZC836A [27] and NTE618 [28], can continuously vary capacitance within this range by adjusting the reverse bias voltage applied to their p-n junctions, thereby enabling arbitrary phase shifts of the reflected waves.
However, the frequencies used for middle- and long-range applications typically employs acoustic frequencies below 50 kHz. Under these conditions, the capacitance of a varactor diode must be varied substantially to effect a desired phase shift in the reflected wave. For example, under the same 50 $\Omega$ impedance-matching conditions used previously, reducing the source frequency to 50 kHz (for mid-range acoustic communication) requires the RIS load capacitance to be tuned from 26 nF to 64 nF in order to shift the reflection phase from $-45^{\circ}$ to $-90^{\circ}$. If the frequency is further reduced to 25 kHz (for long-range communication), the capacitance must be adjusted from 52 nF to 128 nF to obtain the same phase-shift range. Such capacitance variation far exceeds the tuning capability of commercial varactor diodes, which generally have capacitance values below 1 nF, and thus a new method is required to realize arbitrary phase shifts in UA-RIS.
III-B Implementation of Synthetic Reflection
In the proposed UA-RIS, we employ synthetic reflection scheme to generate reflected waves with arbitrary amplitude and phase shift. To implement this scheme, the MCU automatically pairs acoustic reflectors according to the angles of reflected waves. Subsequently, one reflector in each pair produces an in-phase or antiphase component, shifting the phase of the reflected waves by 00° or $180$°. The other reflector generates a quadrature component, shifting the phase of the reflected waves by $90$° relative to the incident signal. By precisely controlling the amplitude of the in-phase, antiphase, and quadrature components, the UA-RIS can synthesize reflected waves with any desired phase in space, as dictated by trigonometric identities:
@@ III-B Implementation of Synthetic Reflection
Figure 3: synthetic reflection for arbitrary phase shift.
Fig. 3 illustrates how synthetic reflection technology operates. In the diagram, each circle represents an acoustic reflector, the size of which is one wavelength to efficiently reflect the incident waves. In the figure, the green plane indicates the direction of the incident wave, while the orange plane designates the target direction for the reflected waves.
To execute synthetic reflection, only reflectors located on the wavefront of the reflected wave (as marked by the black dotted line in the figure) can be paired. For example, reflector 1 can be paired with reflector 2 or 2. With this configuration, the projection positions of the paired reflectors along the direction of the reflected waves align. Consequently, positional differences between paired reflectors do not result in additional phase shifts in the direction of the reflected wave. This setup enables the paired reflectors to function collectively as a single reflector, facilitating precise beam steering.
It is important to note that the phase of the incident wave received by each reflector in a pair may differ. This discrepancy arises because the projection positions of the paired reflectors along the wavefront of the incident signals do not always align, as depicted in Fig. 3, where the green dotted line (representing the wavefront of the incident wave) is not parallel to the black dotted line. Therefore, the UA-RIS must correct these phase differences to ensure that the reflected waves generated by the pair of reflectors remain orthogonal, which is crucial for effective synthetic reflection.
@@
Comparing (3) with (2) reveals that $\left(A_{1}\cos\phi_{1}+A_{2}\sin\phi_{2}\right)$ is the amplitude of the in-phase or antiphase component, while $\left(A_{1}\sin\phi_{1}+A_{2}\cos\phi_{2}\right)$ is the amplitude of the quadrature components for synthetic reflection.
@@
IV Hardware Architecture
This section outlines the hardware design of UA-RIS, beginning with an introduction to the structure of the acoustic reflector and the circuitry designed for efficient synthetic reflection. Following this, we will discuss several critical considerations essential for real-world implementation.
IV-A Structure and Circuit of Acoustic Reflector
IV-A1 Reflector Structure